OpenStudy (anonymous):

Simplify:

jhonyy9 (jhonyy9):

what ?

OpenStudy (anonymous):

That dont make sence

OpenStudy (anonymous):

0.5 + (0.08476/3rdsqrt(D^4/256)) + 1

TheSmartOne (thesmartone):

Can you put it into an equation or atleast provide us with a picture to better assist you.

OpenStudy (anonymous):

yup, one sec.

OpenStudy (anonymous):

I'm sorry, this computer is incredibly slow!!

OpenStudy (anonymous):

@zepdrix How would you do this?

zepdrix (zepdrix):

Hey Mr CornDog :D What do we want to do here? Combine it into a single fraction or something?

zepdrix (zepdrix):

\[\Large\rm 0.5+\frac{0.08476}{\sqrt[3]{\frac{D^4}{256}}}+1\]Simplify this?

OpenStudy (anonymous):

Yes! That's what we want to simplify, into one fraction.

zepdrix (zepdrix):

Combine the 0.5 and 1,\[\Large\rm 1.5+\frac{0.08476}{\sqrt[3]{\frac{D^4}{256}}}\]Apply the cube root to each part separately,\[\Large\rm 1.5+\frac{0.08476}{\frac{\sqrt[3]{D^4}}{\sqrt[3]{256}}}\]Do you remember how to deal with division by a fraction? We can rewrite it as multiplication if we FLIP the fraction in the bottom:\[\Large\rm 1.5+0.08476\cdot \frac{\sqrt[3]{256}}{\sqrt[3]{D^4}}\]

zepdrix (zepdrix):

Bring the 0.08476 into the numerator, and calculate your cube root of 256,\[\Large\rm 1.5+\frac{0.08476\cdot 6.34960}{\sqrt[3]{D^4}}\]

zepdrix (zepdrix):

Then to combine both terms, the 1.5 needs the same denominator as the other term.

zepdrix (zepdrix):

Rewrite the D using rational exponent, \[\Large\rm 1.5+\frac{0.53819}{D^{4/3}}\]Then get that denominator,\[\Large\rm \frac{1.5D^{4/3}}{D^{4/3}}+\frac{0.53819}{D^{4/3}}\quad=\quad \frac{1.5D^{4/3}+0.53819}{D^{4/3}}\quad \color{green}{\checkmark}\]

zepdrix (zepdrix):

Mmm so what do you think? Too many steps? Confusing?

OpenStudy (anonymous):

Very well done, I fully understand. But if we were to multiply that fraction by \[\frac{ 51.65 }{D ^{4} }\] would that be \[\frac{ 77.48D^{\frac{ 4 }{ 3 }}+27.8 }{ D ^{\frac{ 16 }{ 3 }} }\]

OpenStudy (anonymous):

I ask because I'm doing root finding for my numerical methods class, and I was given a pretty ugly function, so I'm trying to simplify it before I start.

zepdrix (zepdrix):

Mmmmm yah that step looks ok! :O

OpenStudy (anonymous):

Thank you very much! It's always appreciated.